OpenStudy (marcelie):

help please determine whether each integral is convergent or divergent. evalute those that are convergent

OpenStudy (marcelie):

\[\int\limits_{-\infty}^{\infty} xe ^{-x^2}\]

OpenStudy (marcelie):

@Astrophysics

OpenStudy (marcelie):

@IrishBoy123

OpenStudy (kainui):

There are several ways to think about this. You can either do a substitution \(x^2=u\) or you can admire that it's an odd function which means the left half is the negative of the right half, so the whole thing cancels out to zero.

OpenStudy (eliesaab):

The suggestion above to use the odd function is not correct. Take for example the divergent series \[ \int_{-\infty}^{\infty} x dx \]

OpenStudy (eliesaab):

I mean divergent integral

OpenStudy (irishboy123):

@eliesaab \(\large \int_{-\infty}^{\infty} x dx\) \(\large = \lim\limits_{t \to \infty} \left[ \dfrac{x^2}{2} \right]_{-t}^{t}\) Whereas @marcelie \( \large \int\limits_{-\infty}^{\infty} x e ^{-x^2} dx \) \( \large = -\dfrac{1}{2} \lim\limits_{t \to \infty} \left [ e ^{-x^2} \right ]_{-t}^{ t }\) @jango_IN_DTOWN posted something uber-similar today @HolsterEmission ?!?! i can't believe that just changing the form of the integral [basically, using limit language] just makes it alright; but this seems to me to be the case.

OpenStudy (eliesaab):

For the definition of improper integral it should be \[\lim_{t->\infty, c->\infty} \left[\frac {x^2}2\right]^t_{-c} \] The above limit does not exist.

OpenStudy (eliesaab):

You cannot restrict only to the case t=c

OpenStudy (eliesaab):

@IrishBoy123

OpenStudy (marcelie):

i think i figured it out lol

OpenStudy (holsteremission):

I agree with @eliesaab - in general, the argument that the function is odd is not enough to determine that the integral converges. You also need the fact that \(f(x)\to0\) as \(x\to\pm\infty\) (and \(f\) must be at least piecewise continuous, I would imagine), similar to how \(\sum\limits_na_n\) has the necessary (but not sufficient) condition \(a_n\to0\) in order to converge.

OpenStudy (irishboy123):

tuvm